OpenStudy (anonymous):

Solve. 3 1 ---- = ------- x x-2

hero (hero):

3(x-2) = x 3x - 6 = x 3x - x = 6 2x = 6 x = 6/2 x = 3

OpenStudy (anonymous):

thanx